# ChemMLLM: Chemical Multimodal Large Language Model

### v1

@@ Appendix I Ablation Study
data augmentation VQGAN MW Pearson ($\uparrow$) LogP Pearson ($\uparrow$) TPSA Pearson ($\uparrow$) Hbd Pearson ($\uparrow$) Hba Pearson ($\uparrow$) Rb Pearson ($\uparrow$) QED Pearson ($\uparrow$)
$\checkmark$ $\checkmark$ 0.71 0.42 0.71 0.45 0.66 0.62 0.34
$\times$ $\checkmark$ 0.46 0.04 0.40 0.08 0.53 0.17 0.26
$\times$ $\times$ 0.55 0.06 -0.05 0.06 -0.06 0.31 0.35
Table 18: Ablation study.
We conduct ablation study on property2img task with a focus on evaluating the impact of Mol-VQGAN training and data augmentation. The effectiveness of the generated molecular images is assessed through the Pearson correlation coefficients between the predicted and ground truth values of several molecular properties, including MW, LogP, TPSA, Hbd, Hba, Rb, and QED. we use a subset of 200 samples of he property2img task and do 5-shot evaluation.
The result is shown in Table 18. When both fine-tuned VQGAN and data augmentation are employed, the model achieves the highest correlation scores across all evaluated properties. Notably, the correlations for MW (0.71), TPSA (0.71), Hba (0.66), and Rb (0.62) indicate that the generated images capture molecular structure and features that align well with the original textual descriptions. This demonstrates the efficacy of our approach in enhancing the semantic fidelity and chemical relevance of the generated visual representations. Removing data augmentation while retaining the fine-tuned VQGAN leads to a significant drop in performance, especially for properties such as LogP (reduced to 0.04) and TPSA (0.40), highlighting the importance of data augmentation in improving the model’s generalization and robustness during training. The performance further deteriorates when both fine-tuning and data augmentation are removed. In this setting, the model yields the lowest correlations, with some properties (e.g., TPSA at -0.05 and Hba at -0.06) exhibiting negative correlation, suggesting that the general VQGAN trained on natural images fails to preserve critical molecular features necessary for reliable property prediction.
In summary, these results clearly demonstrate that both fine-tuning Mol-VQGAN and applying data augmentation play complementary and crucial roles in enhancing the quality and accuracy of chemical multimodal tasks.

### v2

@@ Appendix I Ablation Study
Data Aug VQGAN LLM SFT MW Pearson ($\uparrow$) LogP Pearson ($\uparrow$) TPSA Pearson ($\uparrow$) Hbd Pearson ($\uparrow$) Hba Pearson ($\uparrow$) Rb Pearson ($\uparrow$) QED Pearson ($\uparrow$)
$\checkmark$ $\checkmark$ $\checkmark$ 0.71 ±0.05$\pm$ 0.05± 0.05 0.42 ±0.05$\pm$ 0.05± 0.05 0.71 ±0.05$\pm$ 0.05± 0.05 0.45 ±0.08$\pm$ 0.08± 0.08 0.66 ±0.06$\pm$ 0.06± 0.06 0.62 ±0.05$\pm$ 0.05± 0.05 0.34 ±0.08$\pm$ 0.08± 0.08
$\times$ $\checkmark$ $\checkmark$ 0.46 ±0.13$\pm$ 0.13± 0.13 0.04 ±0.17$\pm$ 0.17± 0.17 0.40 ±0.10$\pm$ 0.10± 0.10 0.08 ±0.07$\pm$ 0.07± 0.07 0.53 ±0.10$\pm$ 0.10± 0.10 0.17±0.07$\pm$ 0.07± 0.07 0.26 ±0.07$\pm$ 0.07± 0.07
$\checkmark$ $\times$ $\checkmark$ 0.24 ±0.16$\pm$ 0.16± 0.16 0.19 ±0.07$\pm$ 0.07± 0.07 -0.02±0.08$\pm$ 0.08± 0.08 0.007 ±0.06$\pm$ 0.06± 0.06 -0.06 ±0.06$\pm$ 0.06± 0.06 0.23 ±0.17$\pm$ 0.17± 0.17 0.36 ±0.07$\pm$ 0.07± 0.07
$\checkmark$ $\checkmark$ $\times$ -0.05 ±0.08$\pm$ 0.08± 0.08 -0.04 ±0.06$\pm$ 0.06± 0.06 -0.01 ±0.04$\pm$ 0.04± 0.04 0.0 ±0.0$\pm$ 0.0± 0.0 0.05 ±0.05$\pm$ 0.05± 0.05 0.02 ±0.07$\pm$ 0.07± 0.07 0.06 ±0.07$\pm$ 0.07± 0.07
Table 19: Ablation study. Data augmentation is written as Data Aug.
We conduct ablation study on property2img task with a focus on evaluating the impact of Mol-VQGAN training, ChemMLLM SFT and data augmentation. The effectiveness of the generated molecular images is assessed through the Pearson correlation coefficients between the predicted and ground truth values of several molecular properties, including MW, LogP, TPSA, Hbd, Hba, Rb, and QED. we use a subset of 200 samples of the property2img task and do 5-shot evaluation.
The result is shown in Table 19. When all Mol-VQGAN training, ChemMLLM SFT and data augmentation are employed, the model achieves the highest correlation scores across all evaluated properties. Notably, the correlations for MW (0.71), TPSA (0.71), Hba (0.66), and Rb (0.62) indicate that the generated images capture molecular structure and features that align well with the original textual descriptions. This demonstrates the efficacy of our approach in enhancing the semantic fidelity and chemical relevance of the generated visual representations. Removing data augmentation while keeping VQGAN and LLM fine-tuning causes substantial performance degradation, most notably for LogP (0.04) and Hbd (0.08), underscoring data augmentation’s critical role in learning meaningful molecular representations. When VQGAN is excluded while maintaining data augmentation and LLM fine-tuning, the model shows inconsistent performance - maintaining moderate correlations for QED (0.36) but failing completely on TPSA (-0.02) and Hba (-0.06), indicating VQGAN’s importance for spatial feature learning. Removing LLM fine-tuning while keeping other components results in near-random performance across all properties (with correlations close to zero), demonstrating that the LLM’s chemical understanding is essential for meaningful molecule image generation.
In summary, these results clearly demonstrate that both two-stage training and data augmentation play complementary and crucial roles in enhancing the quality and accuracy of chemical multimodal tasks.
Appendix J Analysis of hyper-parameters
J.1 Analysis of $\lambda$
[…]
Figure 18: Results on img2caption task with different $\lambda$
Appendix K Limitations
Appendix L Additional Experiment Result
L.1 Result For img2caption
The results for img2caption are shown in Table 16.
L.2 Result For img2property
The results for img2property are shown in Table 14 and Table 15.